Grade 11Mechanics

A man is swimming in a lake in a direction of 30° East of North with a speed of 5 km/hr and a cyclist isgoing on a road along the lake shore towards East at a speed of 10 km/hr. In what direction and withwhat speed would the man appear to swim to the cyclist.

Aditya Maurya
9 Years agoGrade 11

2 Answers

Vikas TU
9 Years ago
Speed of Man = 5 km/hr
Speed of cyclist = 10 km/hr
Angle = 30°
Relative speed of man with cyclist
= √(102+ 52+2*10*5* cos 60
  • √100+25+50
  • √175 = 5√7 m/sec
And the direction of man from the cyclist
  • Ѳ =( 5 sin 60)/(10+5cos60)
           = (5√3/2)/ (10+5/2)
          = (2.5√3)/12.5
           = √3/5
          = 1.732/5 = 0.346
Ѳ = tan-1(0.346)
    = 19.1o
 
Nishtha
7 Years ago
Speed of man =5km/hr
Speed of cyclist=100km/hr
By using vectors
Angle with the horizontal is 120 degree
By applying relative motion
Speed of man w.r. t cyclist=
√(100+25-2*10*5cos60)=5√3 m/sec